OpenStudy (anonymous):

1/2x ^{2}-1/4x-1/4=0 solve.

OpenStudy (chaise):

This is a quadratic equation. a = 1/2 b = 1/4 c = -1/4 \[x=(-b +/- \sqrt{b^2-4ac}/{2a}\] Substitute in your values. \[x=-2\] \[x=1\]

OpenStudy (gw2011):

(1/2)(x^2)-(1/4)(x)-1/4=0 Multiply by 4 to eliminate the denominator 2x^2-x-1=0 (2x+1)(x-1)=0 2x+1=0 2x=-1 x=-1/2 x-1=0 x=1 Therefore, x=-1/2 and x=1